# To See a World in a Spark of Neuron: Disentangling Multi-task Interference for Training-free Model Merging

### v2

@@ 5.4 Merging Vision Models
We also examined the modality of vision by adhering to the experimental setup outlined by Ilharco et al. (2022, 2023). Specifically, we adopted two variants of the CLIP model (Radford et al., 2021), ViT-B/32 and ViT-L/14, as visual encoders (Dosovitskiy et al., 2021). During the fine-tuning process, we optimized the visual encoder across eight distinct tasks while maintaining a fixed text encoder configuration. This comprehensive evaluation spans multiple classification domains, encompassing remote sensing, traffic analysis, and satellite imagery recognition, with evaluations conducted on standard benchmark datasets, including Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), RESISC45 (Cheng et al., 2017), SUN397 (Xiao et al., 2016), and SVHN (Netzer et al., 2011). Table 4 presented the results of NeuroMerging, demonstrating its competitive performance across different validation scenarios. When employing validation data, our method achieved performance improvements of $0.2\%$ for ViT-B/32 and $0.8\%$ for ViT-L/14 over state-of-the-art baselines. In the absence of additional validation, NeuroMerging further improved upon the strongest baseline by $0.5\%$ and $1.0\%$ for ViT-B/32 and ViT-L/14, respectively. These results substantiated the broad model compatibility of our approach.
6 Additional Results and Analysis
6.1 Merging without Validation Sets
When validation data is unavailable, we examine the parameters $\lambda_{1}$ and $\lambda_{2}$ selected according to Section 3.4, where $\lambda_{1}$ is set to zero, as the corresponding subspace is observed to have little impact in Section 6.3. The value of $\lambda_{2}$ is computed based on the L1-norm of masked and unmasked task vectors. Figure 8 and Table 2 and Appendix Tables 2,  13, and 14 present the evaluation of NeuroMerging on NLP and CV tasks across various model sizes, comparing it with existing methods. NeuroMerging achieves the highest average accuracy across all tasks. This demonstrates that our proposed method, with a simple rescaling, outperforms existing methods on average. Specifically, it achieved a 1.4% and 0.7% improvement over the strongest baseline for T5-Base and T5-Large, respectively. For vision models, it outperforms the strongest baseline by 0.5% and 1.0% for ViT-B/32 and ViT-L/14, respectively.
6.2 Role of Neuronal Subspaces
Figure 2 and Appendix Table 17 illustrated the impacts of neuronal subspace decomposition on T5-Large. To examine the impact of each subspace, ablation was performed separately on each subspace by retaining one while removing the other. Retaining the orthogonal subspace while removing the parallel subspace preserved near-perfect performance of finetuned checkpoints or even improved them across most tasks for T5-Large, achieving 88.0% in-domain, 53.9% out-of-domain, and an average of 58.8%. In contrast, keeping the parallel subspace while removing the orthogonal subspace resulted in a significant performance drop.
@@ 6.2 Role of Neuronal Subspaces
Table 5: In/Out-domain Performance of T5-large.
@@ 6.3 Robustness of Hyperparameters
Relationship Between $\lambda_{1}$ and $\lambda_{2}.$ To examine the effects of $\lambda_{1}$ and $\lambda_{2}$, we conducted a grid search with $\lambda_{1}\in[0,1.0]$ and $\lambda_{2}\in[3.0,4.0]$ at $0.1$ intervals, fixing $r=10\%$. As visualized in Figure 5, the performance exhibited column-wise uniformity in the heatmap, indicating insensitivity to variations in $\lambda_{1}$, which aligns with our earlier discussion on the role of the orthogonal subspace in Section 6.2. The optimal performance occurred at $\lambda_{2}=3.6$, attributed to the substantial proportion of masked variables.
Figure 5: Impacts of $\lambda_{1}$ and $\lambda_{2}$ on T5-Large.
@@ 7 Conclusion
In this paper, we revisited model merging from the core principle neuron connectivity and activation that underpin the learning process of recent deep neural networks and LLMs. Specifically, we presented the first exploration into the roles of neuronal mechanisms in the model merging process, by decomposing of task-specific representations into two complementary neuronal subspaces for characterization. It is mathematically shown that neuronal subspaces regulate input sensitivity and task adaptability. Based on these insights, we proposed NeuroMerging, a novel framework designed to reduce task interference within neurons. Our empirical evaluations demonstrate that NeuroMerging achieves superior performance compared to existing approaches on multi-task benchmarks across both vision and natural language processing tasks.
Future work could extend NeuroMerging to larger models or multimodal architectures with more tasks. While our study introduces a neuronal mechanistic perspective for in model merging, it focuses only on neuron connectivity and activation, without considering higher-order neuronal dynamics (e.g., network level connectivity and dynamics). Further research is needed to investigate these factors and deepen our understanding of neuronal mechanisms in model merging, or even in multi-task learning.
@@
While our work provides the first neuronal mechanistic perspective on multi-task interference when merging large models, (1) it remains a partial view of neuronal mechanisms as it does not yet explore higher order neuronal dynamics during merging and inference, which require further investigation. Moreover, this work shares similar limitations with current SOTA model merging methods, including (2) the effectiveness of task arithmetic in model merging relies on selecting fine-tuned checkpoints that are beneficial for specific domains, ensuring they originate from the same pretrained model, and addressing hyperparameter sensitivity; and (3) More effort is needed to develop a mathematical understanding of why and when task arithmetic in model merging works well, despite its simplicity and efficiency.
@@ Appendix A Additional Details
Input: Task-specific models $\tau_{1},\tau_{2},\dots,\tau_{T}$, pretrained model $\theta_{0}$, Mask ratio $r$
@@
A.1 Checkpoints for Llama
Algorithm 1 is the pseudo-code for NeuroMerging.
A.2 Masking
As entries of the task vector with larger magnitudes are more relevant to task-specific adaptation, only the top $r\%$ of $\mathbf{\tau}_{t}$ with the largest magnitudes are kept, while the others are set to zero Yadav et al. (2023). The masked task vector is defined as $\tau^{\text{masked}}_{t}=m_{t}\circ\tau_{t}$, where $m_{t}$ is the mask that keeps the top $r\%$ of the elements of each task vector. In the main manuscript, we use $\tau_{t}$ to represent the masked task vector for readability.
@@ Appendix B Why Merging at the Neuronal Level?
From the statistical analysis of neuronal versus non-neuronal parameters, it is observed that neuronal parameters dominate the T5-Base and T5-Large, showed in Figure. 6. As a result, they play a major role in shaping the model’s learning dynamics, task adaptability, and overall performance. This dominance suggests that understanding and optimizing neuronal parameter interactions is crucial for improving model merging, reducing task interference, and enhancing generalization across diverse tasks.
@@
Method Scale Granularity Level
Fisher Merging[NeurIPS22] Fisher Matrix Parameter
RegMean[ICLR23] Inner Product Matrix Parameter
Task Arithmetic[ICLR23] Uniformed Task
Ties-Merging[NeurIPS23] Uniformed Parameter
DARE[ICML24] $1/(1-p)$ Parameter
LoraHub[COLM24] Evolver Searched Task
AdaMerging[ICLR24] Unsupervised Optimized Layer
PCB-Merging[NeurIPS24] Balancing Matrix Parameter
NeuroMerging (Ours) L1-Norm Neuron
Table 6: Model merging with different scales and granularity levels.
@@ Appendix C Additional Results
C.1 Ablation Study on Merging Functions
Method Avg. Acc
elect + mean 82.2
elect + sum 79.6
mean 79.7
sum 79.8
Table 7: Comparison of $\psi(\cdot)$ on Avg. Accuracy.
C.2 Mask Ratio vs. Performance
When validation is available, we analyze the impact of the mask ratio on performance, showed in Figure 7. Maintaining $r$ at 10–20% improves performance, whereas exceeding this range often leads to degradation. This suggests that an optimal masking ratio balances information retention and redundancy reduction.
Figure 7: Impacts on mask ratio $r$.
Masking Ratio $r$. We observed robust performance across different ratios, with accuracy peaks at $15\%$ and $10\%$ for T5-Base and T5-Large, respectively. Performance variations remain bounded (within $2.5\%$ for T5-Base and $4\%$ for T5-Large) and stabilize beyond $35\%$.
@@ C.1 Ablation Study on Merging Functions
Table 8: NLP Task Accuracy under Orthogonal and Parallel Projections across T5-Large and T5-Base Models
@@
AdaMerging[ICLR2024] 71.1 71.8 48.9 84.9 85.0 65.1
AdaMerging++[ICLR2024] 73.7 74.5 49.2 87.3 87.5 65.6
@@
Table 9: Vision Task Accuracy under Orthogonal and Parallel Projections across ViT-B/32 and ViT-L/14 Models
@@
Method($\downarrow$) Time (s,$\downarrow$) Mem (GB,$\downarrow$) Acc. (%$\uparrow$) Time (s,$\downarrow$) Mem (GB,$\downarrow$) Acc. (%$\uparrow$)
@@
Table 10: Comparison on Computational Efficiency and Accuracy for NLP Tasks
@@
Method($\downarrow$) Time (s,$\downarrow$) Mem (GB,$\downarrow$) Acc. (%$\uparrow$) Time (s,$\downarrow$) Mem (GB,$\downarrow$) Acc. (%$\uparrow$)
@@
Table 11: Comparison on Computational Efficiency and Accuracy for Vision Tasks
C.3 Comprehensive Task-Level Results
We provide all task-level results in T5-Base, T5-Large (Raffel et al., 2020), LLaMA2 (Touvron et al., 2023), ViT-B/32, and ViT-L/14 (Dosovitskiy et al., 2021), respectively. The task-level results of the in-domain experiments for all models can be found in Tables 121314. The task-level results of the out-domain experiments for T5-Base and T5-Large can be found in Tables 15 and 16. Lastly, Table 17 shows the task-level results from Section 6.2 when only one of the neuronal subspace is retained.
C.4 Existing Model Merging Methods in Neuronal Subspaces
To empirically validate the neuronal subspaces, we extended our experiments to various existing model merging methods. As shown in Tables 8 and 9, experimental results demonstrate that applying these merging methods within our proposed orthogonal neuronal subspaces consistently enhances their performance. Notably, this improvement occurs despite these existing methods not explicitly incorporating neuronal mechanisms into their designs. This observation further underscores the significance and motivation of our proposed approach.
C.5 Scalability Evaluation
We further evaluated the scalability of NeuroMerging from the perspective of computational efficiency. As summarized in table 10 and 11 (runtime in seconds, memory usage in GB; lower is better), NeuroMerging consistently achieves superior efficiency compared to existing methods across representative NLP and CV benchmarks. Specifically, it demonstrates comparable or reduced memory usage and improved accuracy relative to TIES-Merging Yadav et al. (2023) ($+2.7\%$ on T5-Large, $+2.3\%$ on ViT-L/14, $+1.7\%$ on LLaMA2), significantly lower runtime and memory overhead compared to PCB-Merging Du et al. (2024), and vastly reduced computational requirements relative to AdaMerging Yang et al. (2024b), which incurs up to $100$ times longer runtime and $10$ times higher memory usage on ViT-L/14 due to additional training steps. In general, unlike prior approaches Yadav et al. (2023); Du et al. (2024); Yang et al. (2024b); Jung et al. (2024); Yang et al. (2024a); Huang et al. (2024b); Lu et al. (2024); Matena and Raffel (2022), which either require loading all parameters simultaneously or introduce costly auxiliary components, NeuroMerging leverages neuronal task vectors aligned with neuronal mechanisms, ensuring scalable, lightweight, and efficient merging.

### v4

@@ 5.4 Merging Vision Models
We also examined the modality of vision by adhering to the experimental setup outlined by Ilharco et al. (2022, 2023). Specifically, we adopted two variants of the CLIP model (Radford et al., 2021), ViT-B/32 and ViT-L/14, as visual encoders (Dosovitskiy et al., 2021). For a fair comparison, we obtained fine-tuned checkpoints from Ilharco et al. (2023), which were consistently utilized across all baseline methods. This comprehensive evaluation spans multiple classification domains, encompassing remote sensing, traffic analysis, and satellite imagery recognition, with evaluations conducted on standard benchmark datasets, including Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2018), GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), RESISC45 (Cheng et al., 2017), SUN397 (Xiao et al., 2016), and SVHN (Netzer et al., 2011). Table 5 presented the results of NeuroMerging, demonstrating its competitive performance across different validation scenarios. When employing validation data, our method achieved performance improvements of $0.2\%$ for ViT-B/32 and $0.8\%$ for ViT-L/14 over state-of-the-art baselines. In the absence of additional validation, NeuroMerging further improved upon the strongest baseline by $0.5\%$ and $1.0\%$ for ViT-B/32 and ViT-L/14, respectively. These results substantiated the broad model compatibility of our approach. For comprehensive results across all tasks and model variants, see Appendix C.1 Tables 14 and 15, with detailed error analysis provided in Appendix C.4.
@@
Table 6: Role of orthogonal and parallel subspaces on T5-Large, details provided in Appendix Table 18.
6 Additional Results and Analysis
6.1 Merging without Validation Sets
When validation data is unavailable, we examined the parameters $\lambda_{1}$ and $\lambda_{2}$ selected according to Section 3.4, where $\lambda_{1}$ was set to zero, as the corresponding subspace was observed to have little impact in Section 6.3. The value of $\lambda_{2}$ was computed based on the L1-norm of masked and unmasked task vectors. Figure 8 and Table 3 and Appendix Tables 3,  14, and 15 presented the evaluation of NeuroMerging on NLP and CV tasks across various model sizes, comparing it with existing methods. NeuroMerging achieves the highest average accuracy across all tasks. This demonstrated that our proposed method, with a simple rescaling, outperformed existing methods on average. Specifically, it achieved a $1.4\%$ and $0.7\%$ improvement over the strongest baseline for T5-Base and T5-Large, respectively. For vision models, it outperformed the strongest baseline by $0.5\%$ and $1.0\%$ for ViT-B/32 and ViT-L/14, respectively.
6.2 Role of Neuronal Subspaces
Figure 2, Table 6 and Appendix Table 18 illustrated the impacts of neuronal subspace decomposition on T5-Large. To examine the impact of each subspace, ablation was performed separately on each subspace by retaining one while removing the other. Retaining the orthogonal subspace while removing the parallel subspace preserved near-perfect performance of finetuned checkpoints or even improved them across most tasks for T5-Large, achieving $88.0\%$ in-domain, $53.9\%$ out-of-domain, and an average of $58.8\%$. In contrast, keeping the parallel subspace while removing the orthogonal subspace resulted in a significant performance drop.
Figure 5: Impacts of $\lambda_{1}$ and $\lambda_{2}$ on T5-Large.
@@ 6.3 Robustness of Hyperparameters
Relationship Between $\lambda_{1}$ and $\lambda_{2}.$ To examine the effects of $\lambda_{1}$ and $\lambda_{2}$, we conducted a grid search with $\lambda_{1}\in[0,1.0]$ and $\lambda_{2}\in[3.0,4.0]$ at $0.1$ intervals, fixing $r=10\%$. As visualized in Figure 5, the performance exhibited column-wise uniformity in the heatmap, indicating insensitivity to variations in $\lambda_{1}$, which aligned with our earlier discussion on the role of the orthogonal subspace in Section 6.2. This highlights a practical strength of our approach: when computational resources are limited, users can simplify their merging tasks by primarily tuning $\lambda_{2}$, as adjusting $\lambda_{1}$ yields relatively subtle effects. The optimal performance occurred at $\lambda_{2}=3.6$, attributed to the substantial proportion of masked variables.
Figure 6: Impacts on mask ratio $r$.
Masking Ratio $r$. When validation is available, we analyzed the impact of the mask ratio on performance, showed in Figure 6. We observed robust performance across different ratios, with accuracy peaked at $15\%$ and $10\%$ for T5-Base and T5-Large, respectively. Performance variations remained bounded (within $2.5\%$ for T5-Base and $4\%$ for T5-Large) and stabilized beyond $35\%$. Maintaining $r$ at $10\%\text{--}20\%$ improved performance, whereas exceeding this range often led to degradation. This suggested that an optimal masking ratio balances information retention and redundancy reduction.
6.4 Ablation Study on Merging Functions
Method Avg. Acc
elect + mean 82.2
elect + sum 79.6
mean 79.7
sum 79.8
Table 7: Comparison of $\psi(\cdot)$ on average accuracy.
6.5 NeuroMerging + AdaMerging
We adopted task-wise AdaMerging (Yang et al., 2024b) for fair comparison, as the layer-wise variant incurs substantial computational cost due to tuning individual layer- and tasks-specific $\lambda$ values via expensive backpropagation on a model larger than the pre-trained one. Further computational efficiency analysis between different methods is detailed in Appendix C.3.
As shown in Table 8, our method consistently improves the performance of both layer-wise AdaMerging and AdaMerging++ when applied to their publicly shared layer-wise $\lambda$ directly. Notably, we observe a greater improvement when combining AdaMerging with NeuroMerging compared to AdaMerging++, suggesting that the TIES component within AdaMerging++ may inadvertently discard information beneficial to NeuroMerging.
Method ViT-B/32 ViT-L/14
Layer-wise AdaMerging 80.1 90.8
+NeuroMerging 82.3 (+2.2) 91.3 (+0.5)
Layer-wise AdaMerging++ 81.1 91.0
+NeuroMerging 82.5 (+1.4) 91.0 (+<0.1)
Table 8: Layer-wise enhancement on NeuroMerging.
@@ 7 Conclusion
In this paper, we revisited model merging from the core principle neuron connectivity and activation that underpin the learning process of recent deep neural networks and LLMs. Specifically, we presented the first exploration into the roles of neuronal mechanisms in the model merging process, by decomposing of task-specific representations into two complementary neuronal subspaces for characterization. It was mathematically shown that neuronal subspaces regulate input sensitivity and task adaptability. Based on these insights, we proposed NeuroMerging, a novel framework designed to reduce task interference within neurons. Our empirical evaluations demonstrated that NeuroMerging achieved superior performance compared to existing approaches on multi-task benchmarks across both vision and natural language processing tasks.
Future work could extend NeuroMerging to larger models or multimodal architectures with more tasks. While our study introduced a neuronal mechanistic perspective for in model merging, it focused only on neuron connectivity and activation, without considering higher-order neuronal dynamics (e.g., network level connectivity and dynamics). Further research is needed to investigate these factors and deepen our understanding of neuronal mechanisms in model merging, or even in multi-task learning.
Acknowledgements
[…]
@@ Acknowledgements
While our work provided the first neuronal mechanistic perspective on multi-task interference when merging large models, (1) it remains a partial view of neuronal mechanisms as it does not yet explore higher order neuronal dynamics during merging and inference, which require further investigation. Moreover, this work shares similar limitations with current SOTA model merging methods, including (2) the effectiveness of task arithmetic in model merging relies on selecting fine-tuned checkpoints that are beneficial for specific domains, ensuring they originate from the same pretrained model, and addressing hyperparameter sensitivity; and (3) More effort is needed to develop a mathematical understanding of why and when task arithmetic in model merging works well, despite its simplicity and efficiency.
@@ Appendix A Additional Details
A.1 Checkpoints for LLaMA
A.2 Masking Ratio
As entries of the task vector with larger magnitudes are more relevant to task-specific adaptation, only the top $r\%$ of $\mathbf{\tau}_{t}$ with the largest magnitudes are kept, while the others were set to zero Yadav et al. (2023). The masked task vector is defined as $t$, where $m_{t}$ is the mask that keeps the top $r\%$ of the elements of each task vector. We used $\tau_{t}$ to represent the masked task vector for readability.
A.3 Singular Value Decomposition
We employed singular value decomposition (SVD) to construct a meaningful coordinate system within the orthogonal subspace $\mathcal{O}$, defined as the null space of the pre-trained neuron weights. Specifically, since $\mathcal{O}$ inherently lacks an explicit coordinate system, we first vertically stacked the task-specific neuron weight vectors $\tau_{t}^{k}$ into a matrix $D$, and then performed SVD on $D$ as follows:
$D=Q_{k}\Sigma_{k}P_{k}^{\top}$ (8)
Here, $k$ denotes the rank, which was set to be the number of tasks, and $P_{k}$ contains the top right singular vectors, which represent the dominant, i.e., principal, directions of task variability in the orthogonal subspace $\mathcal{O}$. These directions serve as the coordinate axes for the orthogonal subspace.
Next, we projected the task weights onto these axes using $DP_{k}$, and apply the merging function $\psi_{\parallel}(\cdot)$ column-wise to aggregate the projected components into a merged task representation $\zeta^{k}$ in the low-dimensional space. The final reconstruction of the merged weights was given by $\zeta^{k}P_{k}^{\top}$.
A.4 Algorithm
Algorithm 1 is the pseudo-code for NeuroMerging.
Input: Task-specific models $\tau_{1},\tau_{2},\dots,\tau_{T}$, pretrained model $\theta_{0}$, mask $m_{t}$, mask ratio $r$, proj matrix $\mathbf{P}$
@@ Appendix B Why Merging at the Neuronal Level?
From the statistical analysis of neuronal versus non-neuronal parameters, it was observed that neuronal parameters dominate the T5-Base and T5-Large, showed in Figure. 7. As a result, they play a major role in shaping the model’s learning dynamics, task adaptability, and overall performance. This dominance suggests that understanding and optimizing neuronal parameter interactions is crucial for improving model merging, reducing task interference, and enhancing generalization across diverse tasks.
@@ Appendix C Additional Results
C.1 Comprehensive Task-Level Results
We provided all task-level results in T5-Base, T5-Large (Raffel et al., 2020), LLaMA2 (Touvron et al., 2023), ViT-B/32, and ViT-L/14 (Dosovitskiy et al., 2021), respectively. The task-level results of the in-domain experiments for all models can be found in Tables 131415. The task-level results of the out-domain experiments for T5-Base and T5-Large can be found in Tables 16 and 17. Lastly, Table 18 showed the task-level results from Section 6.2 when only one of the neuronal subspace was retained.
C.2 Existing Model Merging Methods in Neuronal Subspaces
To empirically validate the neuronal subspaces, we extended our experiments to various existing model merging methods. As shown in Tables 9 and 10, experimental results demonstrated that applying these merging methods within our proposed orthogonal neuronal subspaces consistently enhanced their performance. Notably, this improvement occurred despite these existing methods do not explicitly incorporating neuronal mechanisms into their designs. This observation further underscored the significance and motivation of our proposed approach.
C.3 Scalability Evaluation
We further evaluated the scalability of NeuroMerging from the perspective of computational efficiency. As summarized in table 11 and 12 (runtime in seconds, memory usage in GB; lower is better), NeuroMerging consistently achieved superior efficiency compared to existing methods across representative NLP and CV benchmarks. Specifically, it demonstrated comparable or reduced memory usage and improved accuracy relative to TIES-Merging Yadav et al. (2023) ($2.7\%$ on T5-Large, $2.3\%$ on ViT-L/14, $1.7\%$ on LLaMA2), significantly lower runtime and memory overhead compared to PCB-Merging Du et al. (2024), and vastly reduced computational requirements relative to AdaMerging Yang et al. (2024b), which incured up to $100$ times longer runtime and $10$ times higher memory usage on ViT-L/14 due to additional training steps. In general, unlike prior approaches Yadav et al. (2023); Du et al. (2024); Yang et al. (2024b); Jung et al. (2024); Yang et al. (2024a); Huang et al. (2024b); Lu et al. (2024); Matena and Raffel (2022), which either required loading all parameters simultaneously or introduce costly auxiliary components, NeuroMerging leveraged neuronal task vectors aligned with neuronal mechanisms, ensuring scalable, lightweight, and efficient merging.
C.4 Error Analysis
This section analyzes individual sub-tasks where NeuroMerging’s performance is more than $3\%$ lower compared to the best-performing baseline. The detailed breakdown covers NLP-series, Vision-series, and LLM-series tasks across various settings, including with and without validation, different model sizes, as well as in-domain and out-of-distribution (OOD) evaluations. For the quantitative analysis, we report both the performance differences and the single-task rankings. For the qualitative analysis, we offer possible explanations to interpret these observed differences.
NLP series. Performance details are summarized in Tables 3, 13, 16, and 17. On Winogrande, T5-Large without validation is behind by $6.8\%$ (ranking 3/5), and T5-Base with validation by $6.6\%$ (ranking 6/6); on Quartz, T5-Base with and without validation lag by $6.3\%$ (ranking 4/6) and $3.3\%$ (ranking 3/5), respectively; on Story Cloze, T5-Base without validation trails by $5.5\%$ (ranking 2/5); for Cosmos QA in the out-of-distribution (OOD) setting, T5-Large is lower by $9.7\%$ (ranking 6/8); and on WiC, T5-Large and T5-Base under OOD conditions show performance gaps of $6.1\%$ (ranking 3/8) and $13.6\%$ (ranking 3/8), respectively.
These tasks typically require deeper semantic reasoning or more fine-grained contextual alignment compared to other tasks. Winogrande (Sakaguchi et al., 2020) and WiC (Pilehvar and Camacho-Collados, 2019) both rely heavily on nuanced lexical disambiguation or pronoun resolution. In these scenarios, the multi-task merging process, especially under limited model capacity or validation scenarios, may inadvertently ignore task-specific weights, diluting fine-grained semantic signals. Quartz (Tafjord et al., 2019) and Cosmos QA (Huang et al., 2019) demand coherent integration of multi-sentence causal chains and background knowledge. However, parameter interference from shorter-text tasks occasionally compresses the reasoning chain, thus restricting deep semantic extraction. Story Cloze (Sharma et al., 2018) emphasizes narrative coherence modeling, where different optimal configuration conflicts among merged tasks can weaken sensitivity to long-range dependencies. Overall, the performance gaps identified mainly arise from representational shifts induced by cross-task distributional differences, coupled with mismatches between model capacity and task complexity. These findings indicate promising directions for future research on model merging.
Vision series. Performance details are summarized in Tables 5, 14 and 15. On SUN397 (Xiao et al., 2016), ViT-B/32 with validation lag by $3.3\%$ (ranking 3/8); and on Cars (Krause et al., 2013), ViT-B/32 with validation is behind by $3.5\%$ (ranking 2/8), with results similar to other merging methods except for Fisher Merging.
Fisher Merging demonstrated strong performance on these datasets, likely because it leverages the Fisher Information Matrix to selectively emphasize task-critical parameters on these two datasets (e.g., SUN397 requires panoramic scene composition, while Cars hinges on sub-class visual micro-signatures) through computing the Fisher Information Matrix using second-order derivatives or approximation through backpropagation. However, the high computational cost of the Fisher Information Matrix and its dependence on data make it less practical. As a result, recent model merging methods often avoid using the Fisher Information Matrix despite its effectiveness on these two datasets.
LLM series. Performance details are summarized in Table 4. NeuroMerging does not exhibit any sub-task with a performance drop exceeding $3\%$ in this scenario.
@@ C.4 Error Analysis
Table 9: NLP tasks’ accuracy under orthogonal and parallel projections across T5-Large and T5-Base models.
@@
Task-wise AdaMerging[ICLR2024] 71.1 71.8 48.9 84.9 85.0 65.1
Task-wise AdaMerging++[ICLR2024] 73.7 74.5 49.2 87.3 87.5 65.6
Layer-wise AdaMerging[ICLR2024] 80.1 80.4 50.4 90.8 91.0 66.1
Layer-wise AdaMerging++[ICLR2024] 81.1 81.1 50.3 91.0 91.2 65.9
@@
Table 10: Vision tasks’ accuracy under orthogonal and parallel projections across ViT-B/32 and ViT-L/14 models.
@@
[…]
@@
Table 11: Comparison on computational efficiency and accuracy for NLP tasks.
@@
[…]
@@
Table 12: Comparison on computational efficiency and accuracy for vision tasks.